# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v3

@@ Appendix C Analysis of Performance and Pathological Indications
This section displays the reconstruction errors for several of the networks tested. Normalised and Isotropic networks tend to outperform all other networks, though several exceptions do occur. The results should be taken as preliminary, since there were fewer CIFAR reconstructions computed using unnormalised datasets, due to limited compute resources and performance not being the primary study of this paper — particularly since Isotropic-Tanh is considered a placeholder activation function which is likely suboptimal and only made to appear superficially similar to standard-Tanh Bird [2025a]. Nevertheless, they are indicative of a general trend where isotropic networks tend to outperform anisotropic networks on this specific task.
@@
Notably, in every example, the plots for unnormalised anisotropic networks featured significantly greater noise in reconstruction error on the testing set than the unnormalised isotropic networks. This may be indicative of smoother training dynamics, indicated for isotropic functions Bird [2025a]. However, this observation is insufficiently rigorous to draw a conclusive statement regarding smoother optimisation dynamics in isotropic networks.
@@
Overall, this shows a clear tendency for isotropic-defined functions to outperform anisotropic counterparts across most tests, with only limited exceptions. This is interesting, as the isotropic functions are considered only a placeholder, intended to imitate the appearance of the anisotropic counterparts superficially. Thus, they have not undergone a selection process to optimise their functionality, leveraging their distinctly different isotropic properties, so they can likely be improved substantially. These results are a promising indication, but remain preliminary, and are so far only applicable to autoencoding reconstruction, where one may expect isotropy to be particularly favourable in preserving information for reconstruction.
These improvements may not extend to larger models such as transformers. For example, the structure of these models may also be partially a product of selection over anisotropic primitives, and may constitute a circular reinforcement of such forms Bird [2025a]. Therefore, these improvements for isotropic primitives may not be applicable. A careful approach, and perhaps symmetry-principled redesign/rediscovery, of larger models may be necessary for determining isotropy improvements. Nevertheless, these preliminary results offer some support for their adoption in autoencoding reconstruction tasks, even for these illustrative placeholder activation functions, which were not intended for actual usage Bird [2025a].
Appendix D Further Details: Network Architectures and Training
@@
Autoencoder networks are studied in these results, due to their lack of one-hot classification head, which could impose anisotropy on results [^8] Some anisotropy may remain due to the spatialised nature of the image data; however, this is inherent to the data and the nature of visual datasets. Residual anisotropy through data is not the remit of the hypothesis being tested nor the overarching algebraic symmetry framework being studied. Though its effect was considered and discussed in App. A.. The autoencoders are comprised of an encoding, decoding, and hidden section. Separately categorising the ‘hidden’ layer is perhaps unorthodox but important for the study. All networks tested are derived from the one shown in Fig. 37. This diagrammatic depiction is consistent with open-sourced standardisation available Bird [2025b] at https://github.com/GeorgeBird1/Diagramatic-Neural-Networks.
Figure 37: Shows an autoencoding architecture for reconstruction of CIFAR data, in a standardised diagrammatic key system Bird [2025b]. It is comprised of three blocks: encoder, hidden layers and decoder. The encoder flattens the tensor representations into a vector of dimension 3072 and then performs an affine transformation down to n dimensions. Following this, each hidden layer consists of the application of an activation function, which is either standard or isotropic-tanh, followed by another affine transformation from n to n dimensions. This hidden layer process is repeated $L$ times. Finally, the decoder applies the chosen activation function again, followed by an affine map from n to 3072 dimensions, before reshaping the vector into its classical representation. Various latent layers are drawn from the model, particularly after the encoder, denoted as layer 0, or after a hidden-layer block, denoted as $k$. Therefore, all latent layers of the study are $n$-dimensional. The parameters of all affine maps are optimisable through gradient descent.
This depicts the architectures used across all results, where $n$ varies in width and $L$ varies in depth of the autoencoder model. The latent layer of study is indexed by integers $k\in\left[0,L\right]$. For MNIST LeCun et al. [2010] models, the tensor representations are changed from $\left[32,32,1\right]\rightarrow\left[28,28,1\right]$ and $\left[3072\right]\rightarrow\left[784\right]$ in all instances, with corresponding changes in affine maps. This ensures these results follow from a similar and comparable model.
@@
Throughout all experiments, a batch size of $24$ was maintained, alongside a learning rate for PyTorch’s Paszke et al. [2019] momentum gradient descent, with momentum $0.9$, a learning rate $0.08$, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal Glorot and Bengio [2010] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances Bird [2025a]. However, standard multivariate normal or orthogonal initialisation would have been equally valid choices. Still, Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions Bird [2025a]. All values were kept in float64.

### v4

@@ Appendix D Analysis of Performance and Pathological Indications
This section displays the reconstruction errors for several of the networks tested. Normalised and isotropic networks tend to outperform all other networks, though several exceptions do occur. The results should be taken as preliminary, since there were fewer CIFAR reconstructions computed using unnormalised datasets, due to limited compute resources and performance not being the primary study of this paper — particularly since Isotropic-Tanh is considered a placeholder activation function which is likely suboptimal and only made to appear superficially similar to standard-Tanh [4]. Nevertheless, they indicate a general trend in which isotropic networks tend to outperform anisotropic networks on this task.
These results are of particular interest as isotropic-tanh regularly outperforms conventional tanh. This is despite the preexisting ecosystem of support implicitly developed over the prior eighty years for anisotropic functions, covering function selection, architectural developments, and hyperparameter tuning and optimisation processes. These isotropic functions appear to outperform on these tasks, despite being entirely unoptimised placeholders, not undergoing any form of search, merely superficially resembling tanh for the purpose of this ablation, and lacking the broader ecosystem to support them. Although preliminary, it is felt that this is both promising and intriguing, and supports the development of better-suited isotropic functions which may yield further improved performance.
@@
Notably, in every example, the plots for unnormalised anisotropic networks featured significantly greater noise in reconstruction error on the testing set than the unnormalised isotropic networks. This may be indicative of smoother training dynamics, indicated for isotropic functions [4]. However, this observation is insufficiently rigorous to draw a conclusive statement regarding smoother optimisation dynamics in isotropic networks.
@@
Overall, this shows a clear tendency for isotropic-defined functions to outperform anisotropic counterparts across most tests, with only limited exceptions. This is greatly interesting, as the isotropic functions are considered only a placeholder, intended to imitate the appearance of the anisotropic counterparts superficially. Thus, they have not undergone a selection process to optimise their functionality, leveraging their distinctly different isotropic properties, so they can likely be improved substantially. Moreover, they are not reliant on an ecosystem of support, which has long developed implicitly around their anisotropic counterparts. These results are a promising indication, but remain preliminary, and are so far only applicable to autoencoding reconstruction, where one may expect isotropy to be particularly favourable in preserving information for reconstruction.
These improvements may not extend to larger models such as transformers. For example, the structure of these models may also be partially a product of selection over anisotropic primitives, and may constitute a circular reinforcement of such forms [4]. Therefore, these improvements for isotropic primitives may not be applicable. A careful approach, and perhaps symmetry-principled redesign/rediscovery, of larger models may be necessary for determining isotropy improvements. Nevertheless, these preliminary results offer some support for their adoption in autoencoding reconstruction tasks, even for these illustrative placeholder activation functions, which were not intended for actual usage [4].
Appendix E Network Architectures and Training
@@
Autoencoder networks are studied in these results, due to their lack of one-hot classification head, which could impose anisotropy on results [^14] Some anisotropy may remain due to the spatialised nature of the image data; however, this is inherent to the data and the nature of visual datasets. Residual anisotropy through data is not the remit of the hypothesis being tested nor the overarching algebraic symmetry framework being studied. Though its effect was considered and discussed in App. A.. The autoencoders are comprised of an encoding, decoding, and hidden section. Separately categorising the ‘hidden’ layer is perhaps unorthodox but important for the study. All networks tested are derived from the one shown in Fig. 37. This diagrammatic depiction is consistent with open-sourced standardisation available at https://github.com/GeorgeBird1/Diagramatic-Neural-Networks [5].
Figure 37: Shows an autoencoding architecture for reconstruction of CIFAR data, in a standardised diagrammatic key system. It is comprised of three blocks: encoder, hidden layers and decoder. The encoder flattens the tensor representations into a vector of dimension 3072 and then performs an affine transformation down to n dimensions. Following this, each hidden layer consists of the application of an activation function, which is either standard or isotropic-tanh, followed by another affine transformation from n to n dimensions. This hidden layer process is repeated $L$ times. Finally, the decoder applies the chosen activation function again, followed by an affine map from n to 3072 dimensions, before reshaping the vector into its classical representation. Various latent layers are drawn from the model, particularly after the encoder, denoted as layer 0, or after a hidden-layer block, denoted as $k$. Therefore, all latent layers of the study are $n$-dimensional. The parameters of all affine maps are optimisable through gradient descent.
This depicts the architectures used across all results, where $n$ varies in width and $L$ varies in depth of the autoencoder model. The width $n$ was kept relatively small, as the hypervolume grows exponentially with the width. This resulted in the relative density of isotropic activations, which appear to more fully occupy this hypervolume, falling exponentially with width. Therefore, this limited the experiments to the arbitrarily chosen $n\leq 32$ to strike a suitable balance between ensuring good visualisations of isotropic activations and exploring varied widths.
The latent layer of study is indexed by integers $k\in\left[0,L\right]$. For MNIST [17] models, the tensor representations are changed from $\left[32,32,1\right]\rightarrow\left[28,28,1\right]$ and $\left[3072\right]\rightarrow\left[784\right]$ in all instances, with corresponding changes in affine maps. This ensures these results follow from a similar and comparable model.
@@
Throughout all experiments, a batch size of $24$ was maintained, alongside a learning rate for PyTorch’s [22] momentum gradient descent, with momentum $0.9$, a learning rate $0.08$, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal [9] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances [4]. However, standard multivariate normal or orthogonal initialisation would have been equally valid choices. Still, Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions [4]. All values were kept in float64.
@@
Appendix F Mathematical Notation Glossary
This section will outline the notations used across equations.
  • Various groups, $S_{n}$, $D_{n}$, $B_{n}$, $\mathrm{O}\left(n\right)$, denote the permutation, even-sign-flip permutation, hyperoctahedral and orthogonal groups respectivly.
  • $\mathbf{P}$ would indicate a matrix representation of a permutation group, similarly $\mathbf{R}$ would indicate a matrix representation for an orthogonal group. These are in standard representations. Mapping a group element to a matrix representation can be denoted by a function $\rho:\mathcal{G}\rightarrow\mathrm{GL}_{n}\left(\mathbb{R}\right)$, which indicates a map from the set of group elements $\mathcal{G}$ to a $n\times n$ invertible matrix that is a member of the real General-Linear, $\mathrm{GL}$, group.
  • Members of sets are denoted $g\in G$, and usual set-notation is assumed, such as subsets $A\subset B$.
  • Multivariate functions such as $\mathbf{f}$, provide a map as follows $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, with dimensionality $n$ and $m$ defined.
  • Standard basis vectors are denoted $\hat{e}_{i}$, where this is the $i^{\text{th}}$ vector of the basis set $\left\{\hat{e}_{i}|\forall i\right\}$ or sometimes denoted $\left\{\hat{e}_{i}\right\}_{\forall i}$ in shorthand.
  • Vector hats, $\hat{x}$, denote unit-normalised vectors, compared to unnormalised general vectors given by $\vec{x}$. Vectors tangent to others may be denoted $\vec{x}_{\perp}$, whilst parallel vectors denoted by $\vec{x}_{\parallel}$.
  • The vector norm is given by $\left\|\vec{x}\right\|_{2}\equiv\left\|\vec{x}\right\|$. This generalises to $l$-norms when explicitly given as $\left\|\vec{x}\right\|_{l}$
  • In general vector spaces may be denoted $\mathbb{R}^{n}$ or with a particular bound $\left[0,1\right]^{n}$, and may be generalised with multiple indices to $\mathbb{R}^{n\times m\times i\times\cdots}$. This extends to spaces such as $\mathcal{S}^{n}$, which is the unit-norm hyper-sphere, where vectors are all unit-normalised.
  • Jacobians are denoted $\mathcal{J}^{\mathbf{f}}$ for their respective function $\mathbf{f}$.
  • The Einstein summation convention is used for sums and indices.
  • The identity matrix is denoted $\mathrm{I}$.
  • Hadamard products are denoted as $\odot$.
  • The operation $\operatorname{diag}$ produces a diagonal matrix, with entries specified by the input to the $\operatorname{diag}$ operation, generally this is given by $\operatorname{diag}\left(\vec{x}\right)=\left(\vec{x}\vec{1}^{T}\right)\odot\mathrm{I}=\left(\vec{1}\vec{x}^{T}\right)\odot\mathrm{I}$ or symmetrised to $\frac{1}{2}\left(\vec{1}\vec{x}^{T}+\vec{x}\vec{1}^{T}\right)\odot\mathrm{I}$, where $\vec{1}$ is the vector of 1s.